# Hyper-Hermitian Weyl Double Copy

### v1

@@ 4.3 Potentials with $\Theta_{0}=0$
Now we consider solutions $\Theta_{A}$ of the form (46) with $a_{1}=0$ and $b_{4}=b_{2}+b_{3}$, thus the potential is
@@
therefore we have that the manifold $\cal{M}$ obtained from (86) is hyper-Hermitian. The corresponding metric obtained from this potential is given by
@@
if we consider eq. (46) with the parameters $a_{4}=-1$ and $b_{4}=b_{2}+b_{3}$, we obtain the same metric. So we only concentrate on these parameters. The same happens for $b_{1}=0$, $a_{4}=a_{2}+a_{3}$ and $b_{4}=-1$, $a_{4}=a_{2}+a_{3}$, which being symmetric parameters give us similar solutions. Now we compute the contracted Weyl curvature, which reads
@@
This metric is type III, then the double copy (17), is satisfied when Maxwell spinors are given by
@@
Notice that is result is consistent with the formula for the generic case (59) by taking the appropriated values of the parameters.
As a special case we consider $b_{2}=1$ and $b_{3}=0$ to get
@@
This solution is hyper-Hermitian and non-hyper-Kähler. Some plots of the electric and magnetic field associated with $\Phi^{(2)}_{AB}$ are shown in Figure 3. The electric and magnetic field was obtained from (95) using the flat coordinates (74).
@@
Table 2: Principal null directions of the examples in double copy. The multiplicity $m$ is next to the null directions. Non-elementary Function is type I and Sparling-Tod is type N.
Figure 3: Electric and magnetic field of subsection (4.3) with $t=0$ and $z=1$. We plotted (95) depending only on coordinates (74).
@@ 4.4 Constant Solution
Figure 4: Electric and magnetic field of subsection (4.4) corresponding to constant solution with $t=z=0$. To obtain electromagnetic fields we plotted eq. (101) in local coordinates (74). The first figure corresponds to the static electric field and in the second one to the static magnetic field.
@@
In Figure 4 we plot the static electric and magnetic fields (101). They correspond to the constant solution with $t=z=0$.
@@
where in the present case $J_{AA^{\prime}}$ is no conserved
@@
This result can be obtained also from eq. (59) taking the appropriated choice of the parameters.
We use here a source in the gauge sector considering that Maxwell’s spinor will not be self-dual. The electric field is constant and is mapped by the double copy because the function $F_{A}$ is constant. The scalar field $\Phi$ is given by
@@
Table 3: In this table is sketched the five examples discussed in Section 4. It contains a classification of the worked metrics, in terms of their Petrov types and whether it corresponds with the hyper-Kähler or hyper-Hermitian types.
@@ 4.5 A Non-elementary Function
Our procedure is not only limited to the case of $\Theta_{A}$ given by eq. (36), but also applies to other potentials that are solutions to the equation (16). In this case, there is a class of solutions of the form [57]:
@@
where $k$ and $l$ are integers and $a,b$ are complex constants. There is a family of solutions depending on the specific values of the complex coefficients $a$ and $b$ and values for the integer power $k$. Hence the metric is given by
@@
From this metric, we calculate Weyl spinor curvature and its contraction with four spinors $\xi$’s
@@
Monomials arising from eq. (24) are given by
@@
The double copy imply
@@
Then Maxwell’s field strength spinors are given by
@@
[…]
@@ 5 Final Remarks
In the present article, we studied the Weyl double copy for hyper-Kähler and hyper-Hermitian manifolds. We used local null coordinates, and we showed that there exists a double copy for hyper-Hermitian or hyper-Kähler manifolds of type I, III and N. These were found for six classes of solutions (see Tables 2 and 3), inspired by the Sparling-Tod. We discussed the double copy for the hyper-Kähler case in Sections 4.1 and 4.2, where there is a class of solutions more general, but for certain values of parameters, it is recovered the usual Sparling-Tod double copy. It was confirmed that for the case of hyper-Kähler manifolds all our solutions satisfied Plebański second heavenly equation. The fields we plotted have asymptotes similar to the Sparling-Tod example, thus we suggest that this is explained by the fact that our solutions were inspired by Sparling-Tod and Eguchi-Hanson metrics. Specifically, our families of solutions contain more singular points than the previously mentioned ones. One generic feature of the hyper-Hermitian case we found is that in general there are two different Maxwell’s field strengths and that one of these fields is associated to a source with a calculable current, which in general is not conserved (see eq. (59) in subsection (3.3)). This is compatible with the fact that in general the hyper-Hermitian spaces are not Ricci-flat. In general terms this current is non-conserved, but in some simpler cases, it can be conserved.
The next interest would be looking for the existence of double copy for solutions of Petrov type II. The origin of double copy suggests the possibility of describing type II in terms of Maxwell’s spinors of the single copy. It would be interesting to study these cases or their explanation of their absence to shed light on the double copy correspondence itself and the one of hyper-Hermitian manifolds. Another open problem that we will address in the near future is the twistor description of the double copy for hyper-Hermitian manifolds. Finally, we would like to seek whether there exists a double copy of Kerr-Schild for hyper-Hermitian manifolds without being hyper-Kähler.
Acknowledgments
It is a pleasure to thank M. Dunajski for useful insights and correspondence. G. Robles would like to thank CONACyT for a grant.

### v2

@@ 4.3 Potentials with $\Theta_{0}=0$
Now we consider solutions $\Theta_{A}$ of the form eq. (46) with $a_{1}=0$ and $b_{4}=b_{2}+b_{3}$, thus the potential is
@@
therefore we have that the manifold $\cal{M}$ obtained from eq. (86) is hyper-Hermitian. The corresponding metric obtained from this potential is given by
@@
if we consider eq. (46) with the parameters $a_{4}=-1$ and $b_{4}=b_{2}+b_{3}$, we obtain the same metric. So, we only concentrate on these parameters. The same happens for $b_{1}=0$, $a_{4}=a_{2}+a_{3}$ and $b_{4}=-1$, $a_{4}=a_{2}+a_{3}$, which being symmetric parameters give us similar solutions. Now we compute the contracted Weyl curvature, which reads
@@
We have a type III metric, then the double copy in eq. (17), is satisfied when Maxwell spinors are given by
@@
Notice that this result is consistent with the formula for the generic case in eq. (59) by taking the appropriate values of the parameters.
As a special case, we consider $b_{2}=1$ and $b_{3}=0$ to get
@@
This solution is hyper-Hermitian and non-hyper-Kähler. Some plots of the electric and magnetic fields associated with $\Phi^{(2)}_{AB}$ are shown in Figure 3. The electric and magnetic fields were obtained from eq. (95) using the flat coordinates in eq. (74).
@@
Table 2: Principal null directions of the examples in double copy. The multiplicity $m$ is next to the null directions. Non-elementary Function is type I metric and Sparling-Tod is type N metric.
Figure 3: Electric field of subsection (4.3) with $t=0$ and $z=1$. We plotted eq. (95) depending only on the coordinates of eq. (74). The magnetic field is proportional to the electric field, and therefore it is not plotted.
@@ 4.4 Constant Solution
(a) Electric field.
(b) Magnetic field.
Figure 4: Electric and magnetic field of subsection (4.4) corresponding to constant solution with $t=z=0$. To obtain electromagnetic fields we plotted eq. (101) in local coordinates of eq. (74).
@@
In Figure 4 we plot the static electric and magnetic fields in eq. (101). They correspond to the constant solution with $t=z=0$.
@@
where in the present case $J_{AA^{\prime}}$ is not conserved
@@
This result can be obtained also from eq. (59) taking the appropriate choice of the parameters.
We use here a source in the gauge sector considering that Maxwell spinor will not be self-dual. The electric field is constant and is mapped by the double copy because the function $F_{A}$ is constant. The scalar field $\Phi$ is given by
@@
Table 3: This table sketches the five examples discussed in Section 4. It contains a classification of the worked metrics, in terms of their Petrov types and whether it corresponds with the hyper-Kähler or hyper-Hermitian types.
@@ 4.5 A Non-elementary Function
Our procedure is not only limited to the case of $\Theta_{A}$ given by eq. (36), but also applies to other potentials that are solutions to the eq. (16). In this case, there is a class of solutions of the form ref. [60]:
@@
where $k$ and $l$ are integers, and $a,b$ are complex constants. There is a family of solutions depending on the specific values of the complex coefficients $a$ and $b$, and values for the integer power $k$. Hence, the metric is given by
@@
From this metric, we calculate the Weyl spinor curvature and its contraction with four spinors $\xi$’s
@@
The monomials arising from eq. (24) are given by
@@
The double copy implies
@@
Then Maxwell field strength spinors are given by
@@
where $c_{3}\in\mathbb{R}$ is a constant and $\Phi$ satisfies the wave equation. We do not need to use eq. (18) because here $\Phi$ is simpler. We then observe that eqs. (110), (126), (129), and (130) satisfy eq. (17), which shows that our procedure is possible for other holomorphic functions.
@@ 5 Final Remarks
In the present article, we studied the Weyl double copy for hyper-Kähler and hyper-Hermitian manifolds. We used local null coordinates, and we showed that there exists a double copy for hyper-Hermitian and hyper-Kähler manifolds of type I, III and N. These were found for six classes of solutions (see Tables 2 and 3), inspired by the Sparling-Tod. We discussed the double copy for the hyper-Kähler case in Sections 4.1 and 4.2, where there is a class of solutions more general, but for certain values of parameters, it is recovered the usual Sparling-Tod double copy. It was confirmed that for the case of hyper-Kähler manifolds all our solutions satisfied Plebański second heavenly equation. The fields we plotted have asymptotes similar to the Sparling-Tod example. Thus we suggest that this is explained by the fact that our solutions were inspired by Sparling-Tod and Eguchi-Hanson metrics. Specifically, our families of solutions contain more singular points than the previously mentioned ones. One generic feature of the hyper-Hermitian case we found is that in general there are two different Maxwell’s field strengths and that one of these fields is associated with a source with a calculable current, which is not conserved (see eq. (59) in subsection (3.3)). This is compatible with the fact that in general the hyper-Hermitian spaces are not Ricci-flat. Regarding this point, it is a known fact that in the single copy of Einstein-Maxwell theory there are two U(1) gauge fields, one of them has a source related to the source current of the ordinary Maxwell field [18, 20, 31]. It would be interesting to study if this fact is linked to our result for the hyper-Hermitian single copy with two Abelian gauge fields.
The next interest would be looking for the existence of double copy for solutions of Petrov type II. The origin of double copy suggests the possibility of describing type II metric in terms of Maxwell spinors of the single copy. It would be interesting to study these cases or their explanation of their absence to shed light on the double copy correspondence itself and the one of hyper-Hermitian manifolds. Another open problem that we will address in the near future is the twistor description of the double copy for hyper-Hermitian manifolds. Finally, we would like to seek whether there exists a double copy of Kerr-Schild for hyper-Hermitian manifolds without being hyper-Kähler.
Acknowledgements
It is a pleasure to thank A. Luna for very useful comments and suggestions and to M. Dunajski for useful insights and correspondence. G. Robles would like to thank CONAHCyT for a grant.